OpenStudy (anonymous):

lim as x approaches pi/4 from sin4x*cos2x/(x-pi/4)^2

hartnn (hartnn):

try the substitution x=pi/4 = y

hartnn (hartnn):

x- pi/4 = y

OpenStudy (anonymous):

and then?

hartnn (hartnn):

find sin 4x and cos2x in terms of y

hartnn (hartnn):

since x= y+pi/4 what will be sin4x =... ?

OpenStudy (anonymous):

should i expand sin4x and cos 2x first or? what should i do?